# Bone: Block-Affine Adaptation of Large Language Models

### v5

@@ Abstract
Low-Rank Adaptation (LoRA) has achieved remarkable training results by freezing the original weights and training only low-rank matrices, establishing itself as the predominant fine-tuning method for LLMs. Many LoRA variants have emerged, yet they lack a design tailored to the characteristics of LLM weights and fail to leverage the original weights effectively. To address the sparsity of LLM weights, and drawing inspiration from GQA and MQA, we propose Block-Affine Adaptation (Bone), a novel PEFT technique distinct from LoRA. By dividing the original weights into multiple subspaces that share a single matrix for weight updates, Bone simplifies the process by requiring the trainable matrix to be initialized to zero, eliminating the need for complex initialization as in some LoRA variants. Compared to LoRA, Bone significantly reduces memory usage and achieves faster computation. Evaluation of both NLU and NLG tasks demonstrates that Bone substantially outperforms LoRA and its variants. Inspired by Pissa, we propose a new theory called “Weight Guide” to better utilize the information embedded in the original weights. This approach extracts valuable information through a linear transformation of the original weight matrix using a trainable matrix. To validate the effectiveness of “Weight Guide” we combined it with Bone to create a new structure called Block-Affine Transformation (Bat), and ablation experiments confirmed the effectiveness of “Weight Guide”.
@@ 1 Introduction
Large models have been integrated into various industries, revolutionizing many traditional technologies Radford et al. (2019); Raffel et al. (2020). However, general-purpose large models often struggle to meet the needs of all downstream tasks, making it necessary to fine-tune base models for specific scenarios. Full-scale fine-tuning of large models is computationally costly; for example, finetuning the LLaMA2-7B Touvron et al. (2023) model with bfloat16 Wang & Kanwar (2019) precision requires around 60GB of VRAM. In contrast, PEFT (Parameter-Efficient Fine-Tuning)Xu et al. (2023) techniques could reduce the VRAM requirement to fit into a 24GB VRAM GPU. As a result, numerous PEFT techniques and quantization methods have emerged to reduce the training costs of large models. LoRA (Low-Rank Adaptation) Hu et al. (2021) has become one of the most popular PEFT methods due to its small tunable parameter size, its effectiveness, and the possibility of zero inference overhead after finetuning.
LoRA significantly reduces memory usage by freezing the original weights $W$ and updating two low-rank matrices $A$ and $B$. Typically, either $A$ or $B$ is initialized to zero, ensuring that the initial state of LoRA is consistent with the pre-trained model, The figure  2 illustrates the structure visualization. However, extensive experiments Ding et al. (2023); Liu et al. (2024b); Biderman et al. (2024) have shown that LoRA’s convergence is significantly slower compared to full fine-tuning. This slow convergence is likely due to the small gradients caused by the zero initialization of either $A$ or $B$. To address this issue, researchers have proposed several LoRA variants, such as LoRA+ Hayou et al. (2024), PISSA Meng et al. (2024), and LoRA-GAWang et al. (2024). Despite their excellent performance, these LoRA variants inevitably introduce complexity into the fine-tuning process and reduce structural flexibility. For example, LoRA+ requires manual adjustment of different learning rates for $A$ and $B$; PISSA necessitates performing SVD decomposition on $W$ at the beginning of training, which can be time-consuming when the model parameters are large. As research on LoRA becomes increasingly saturated, it is essential to explore new directions for PEFT techniques and effectively leverage the unique characteristics of LLMs.
In this work, we first designed Block-Affine Adaptation (Bone), which divides the original weights into multiple subspaces that share a single trainable matrix for updates, making it more efficient than LoRA’s two low-rank matrices. Secondly, to address the limitations of LoRA and fully utilize the information in the original weights, we propose a new theory, “Weight Guide”. “Weight Guide” enables feature extraction by applying a simple linear transformation to the original weights. Finally, to validate the effectiveness of “Weight Guide”, we combined Bone with “Weight Guide” to develop Block-Affine Transformation (Bat).
@@
Our extensive evaluation shows that Bone not only excels in both Natural Language Understanding (NLU) and Natural Language Generation (NLG) tasks but also significantly outperforms LoRA and its variants. Additionally, Bone retains the advantages of LoRA, such as ease of use and no additional computational overhead for the model. In NLG tasks, Bone demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variant Pissa across the board. To verify the feasibility of Bone, we conducted experiments on two different LLM architectures (LLaMA2 Xu et al. (2023), RWKV6 Peng et al. (2024)). As shown in Figure  1, Bone achieves the fastest convergence, and the results on the test set demonstrate its superior data fitting and generalization capabilities ( Table  2). As a completely new structure distinct from LoRA, Bone offers clear improvements in computational efficiency and memory savings, as detailed in table  7.
@@
1. We propose a novel PEFT technique called Block-Affine Adaptation (Bone). Bone outperforms LoRA and its variants across various fine-tuning tasks. Additionally, Bone is more memory-efficient and computationally faster compared to LoRA.
2. To effectively leverage the implicit information in the original weights, we propose the “Weight Guide” theory. By integrating this theory with Bone, we design Block-Affine Transformation (Bat), demonstrating the effectiveness of “Weight Guide”.
(a) llama2-7b
(b) Train Steps
Figure 1: The left image shows the loss curve for LLaMA2-7B fine-tuned on the MetaMathQA dataset, with the first 100 steps highlighted for closer observation. Comparing the loss curves reveals that Bone demonstrates superior fitting ability across various architectures and parameter settings. Additionally, Bone exhibits a rapid decrease in loss within the first 100 steps, highlighting its effectiveness. The left image is a demonstration of the training state.
@@ 2 Related Works
The adapter method does not require fine-tuning all the parameters of the pre-trained model. Instead, it introduces a small number of task-specific parameters to store knowledge related to that task, thereby reducing the computational demands of model fine-tuning. Prefix tuning is a lightweight fine-tuning method for generative tasks. It adds a continuous, task-specific vector sequence, called a prefix, to the input. Unlike prompts, prefixes are entirely composed of free parameters and do not correspond to actual tokens. Compared to traditional fine-tuning, prefix tuning only optimizes a set of prefixes related to specific tasks, without any change to the original model.
Prompt tuning defines a unique prompt for each task, prepending it to the input data while freezing the pre-trained model during training. LoRA allows us to indirectly train certain dense layers in a neural network by optimizing low-rank decomposition matrices that adapt these layers while keeping the pre-trained weights unchanged. This approach significantly addresses the inference latency introduced by adapters and the fitting challenges of prefix tuning.
In LoRA, the adapter matrices $A$ and $B$ are updated with the same learning rate, but using the same rate for both may not effectively learn the features. LoRA+ extended this method by introducing independent learning rates for matrices $A$ and $B$ with a fixed ratio, improving the method’s efficiency. The DoRA Liu et al. (2024b) method combines weight decomposition to achieve learning capabilities similar to full fine-tuning without sacrificing LoRA’s inference efficiency. PiSSA optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction. Using Singular Value Decomposition (SVD), PiSSA initializes the dominant singular values and vectors to train these matrices, while keeping the residual matrix static during fine-tuning. OLoRA Büyükakyüz (2024) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal. This orthogonal initialization helps maintain the stability of the parameter space during optimization. LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $A$ and $B$ by computing the initial gradient, thereby closely approximating full fine-tuning.
@@ 3 Method
3.1 Motivation
The primary motivations are as follows:
Firstly, current research on LoRA has reached a saturation point, underscoring the need for a new direction in Parameter-Efficient Fine-Tuning (PEFT) techniques.
Secondly, existing PEFT methods lack a thorough exploration and effective utilization of the original model weights.
3.2 Bone: Block-Affine update matrices
To design a novel PEFT structure that rivals LoRA, we conducted an in-depth study of dense matrices. Current theories suggest that LLM weights exhibit sparse properties. For example, LoRA utilizes low-rank matrices to update dense matrices, while techniques like GQA Ainslie et al. (2023) and MQA Shazeer (2019) employ shared mechanisms to reduce parameter overhead with remarkable effectiveness.
Inspired by these approaches, we hypothesize that dividing LLM weights into multiple subspaces allows different subspaces to share a single low-rank matrix for updates. As illustrated in the figure  2, we name this structure Block-Affine Adaptation (Bone).
It is evident that the ways to partition LLM weights are diverse, as illustrated in the figure. Different partitioning methods influence the shape of the Bone matrix, the computation logic, and ultimately, the processing speed. In this work, we propose the Best Bone structure. For a pre-trained weight matrix $W\in\mathbb{R}^{d\times k}$, we update it using a single matrix: $W+\Delta W=W+expand(B)$, where $B\in\mathbb{R}^{r\times d}$, and the rank $r\ll k$. This design not only retains the simplicity and ease of use characteristic of LoRA but also achieves higher computational efficiency. The calculation is expressed as follows:
$Y=Wx+\Delta Wx=Wx+sum(x)B$ (1)
(a) Bone instruciton
Figure 2: Bone requires training only a single low-rank matrix initialized to zero. Best Bone represents the optimal design structure of Bone. Subsequent experiments default to using Best Bone.
3.3 “Weight Guide”
It is well-known that LLMs have extremely large parameters, with complex interactions among their internal weights, but most LoRA variants utilize original weight information minimally, with only a few leveraging it during initialization to enhance performance. To increase the utilization of this information and promote internal interactions among weights, we propose a novel concept called “Weight Guide” This approach ensures that the trainable matrices are consistently guided and constrained by the original weights throughout every step of the training process. As illustrated in Figure  1(b), LoRA variants interact with the original weights only once during initialization to extract essential components. In contrast, the Bone structure is guided (constrained) by the original weights after each update, significantly increasing the new weights’ utilization of the original weights. As the training steps increase, the influence of the original weights becomes more pronounced. This is evident from the loss curves, which show that the differences between Bone and other methods grow larger in the later stages of training.
So, how can we extract the critical information from the original weights? A straightforward approach is to apply a learnable matrix directly to the original weights using the Hadamard product. This method essentially assigns a weight to each element, allowing the learnable matrix to autonomously determine the importance of individual elements. Although this method is simple and straightforward, the elements within the weights remain isolated, making it impossible to learn interactions between them. Therefore, we replace the Hadamard product with matrix multiplication, essentially applying a simple linear transformation. This approach links the entire weight matrix space, providing a better chance to discover the optimal solution. The corresponding formula is shown below:
$\Delta W_{n,n}$ $=W_{n,n}\odot bone_{n,n}$ (2)
$\Delta W_{n,n}$ $=W_{n,n}\otimes bone_{n,n}$
@@ 3.4 Bat: Block-Affine Transformation
In the previous section, we introduced Bone, which is both simple and efficient. However, applying the same updates to different subspaces of the weight matrix $W\in\mathbb{R}^{d\times k}$ is clearly counterintuitive. To allow for more flexibility in the update directions of different subspaces, we combine Bone with “Weight Guide”, proposing a new method called Block-Affine Transformation (Bat). To balance computational efficiency, we divide the original weights into smaller subspaces compared to Bone. Here, we name the trainable matrix $bone\in\mathbb{R}^{r\times d}$, where its rank $r$ is replaced with block size $b$. The formula for Bone-col is as follows:
$W_{k/b,d/b,b,b}$ $=Reshape(W_{d,k})$ (3)
$bone_{d/b,b,b}$ $=Reshape(bone_{b,d})$
$\Delta W_{d,k}$ $=Reshape(W_{k/b,d/b,b,b}\otimes bone_{d/b,b,b}+bone_{d/b,b,b})$
Similar to Bone, Bat can also adopt various grouping methods, as illustrated in the figure 3. In Section 6, we compared the performance of different partitioning strategies for the structure. The formula for Bone-row is as follows:
$W_{d/b,k/b,b,b}$ $=Reshape(W_{d,k})$ (4)
$bone_{k/b,b,b}$ $=Reshape(bone_{b,k})$
$\Delta W_{d,k}$ $=Reshape(W_{d/b,k/b,b,b}\otimes bone_{k/b,b,b}+bone_{k/b,b,b})$
(a) Bone-both
(b) Bone-col
(c) Bone-row
[…]

### v6

# DiSHA: Dimension-Sharding Adaptation with Fast Convergence and Fast Computation

@@ Abstract
Low-Rank Adaptation (LoRA) leverages the low intrinsic rank of weight updates in Large Language Models (LLMs), establishing a Parameter-Efficient Fine-Tuning (PEFT) paradigm. However, LoRA suffers from slow convergence. We introduce Dimension-Sharding Adaptation (DiSHA), which expands the PEFT design space to unlock lower intrinsic ranks and faster convergence by default. Within DiSHA’s design space, we propose Block Affine Adaptation (Bone), a computationally efficient structure that delivers both high performance and efficiency. While certain DiSHA configurations may result in colinear updates to weight shards, we address this with Block Affine Transformation Adaptation (BAT), a nonlinear variant of DiSHA. BAT introduces nonlinearity by combining trainable matrices with original weight shards in a nonlinear manner, inducing nonlinearity in matrix updates without introducing additional parameters. Empirical results show that Bone, under the DiSHA framework, consistently outperforms LoRA variants in both NLG and NLU tasks, with significantly improved computational efficiency. Further analysis demonstrates that BAT enhances model capabilities by leveraging its nonlinear design.
@@ 1 Introduction
The emergence of Large Language Models (LLMs) has fundamentally transformed many traditional technologies  Radford et al. (2019); Raffel et al. (2020). However, general-purpose large models often struggle to meet the needs of all downstream tasks, making it necessary to fine-tune base models for specific scenarios. Full-scale fine-tuning of large models is computationally costly. As a result, numerous Parameter-Efficient Fine-Tuning (PEFT)Xu et al. (2023) techniques and quantization methods have emerged to reduce the training costs of large models. Low-Rank Adaptation(LoRA) Hu et al. (2021) has become one of the most popular PEFT methods due to its small tunable parameter size, its effectiveness, and the possibility of zero inference overhead after fine-tuning. The figure  3 illustrates the structure visualization.
32641282560612182430rankgradient normftfrozen ftLoRABigLoRAPiSSADiSHA
Figure 1: the initial gradient norm of model training.
However, extensive experiments Ding et al. (2023); Liu et al. (2024b); Biderman et al. (2024) have shown that LoRA’s convergence is significantly slower compared to full fine-tuning. To address this issue, researchers have proposed several LoRA variants, such as LoRA+ Hayou et al. (2024), PISSA Meng et al. (2024), and LoRA-GAWang et al. (2024). These methods have all brought unexpected improvements. By adopting different initialization strategies to influence the model’s training gradients, they have accelerated LoRA’s convergence speed.
Different initializations of LoRA variants accelerate convergence essentially by increasing the initial gradients during training or aligning them with the full-scale training gradients. However, this also makes the initialization process significantly more complex, leading to increased initialization time overhead. To address the issue of complexity and explore the specific reasons for LoRA’s slow convergence, we conducted extensive experiments on initial gradient testing. Ultimately, we concluded that LoRA’s slow convergence is influenced not only by initialization but also by its inherent structural limitations. A simpler update structure may be more suitable for fine-tuning LLMs.
In this work, we first propose a novel framework called Dimension-Sharding Adaptation (DiSHA) to design PEFT with lower ranks and faster convergence. In DiSHA, the pre-trained weights $W_{0}$ are frozen and divided into multiple shards during training, with all shard weights sharing a trainable matrix of the same size as a shard to perform updates, significantly reducing the number of trainable parameters. The design space of DiSHA is vast, allowing pre-trained weights to be divided into dimensions of arbitrary sizes 2. Computational efficiencies vary among design choices. We proposed an efficient and user-friendly structure named Bone within a subset of the DiSHA design space and validated its effectiveness through extensive experiments. Of course, we also observed that certain subsets of the DiSHA design space exhibit limitations: all shard weight updates are colinear, which could potentially hinder model performance. To address this potential issue, we propose the Bat method. Without introducing additional trainable matrices, Bat uses the trainable matrix to interact with the original weight shards, transforming the linear updates between the shards into non-linear updates. This approach enhances the model’s capability while maintaining efficiency.
Figure 2: partition of weight matrix $W_{0}$.
@@
Our extensive evaluation shows that Bone demonstrates excellent performance on both Natural Language Understanding (NLU) and Natural Language Generation (NLG) tasks, significantly surpassing LoRA and its variants. Moreover, as shown in Figure  5, Bone also achieves a significantly faster convergence rate compared to PiSSA. This demonstrates the feasibility of the DiSHA framework. Furthermore, Bone’s unique design not only resolves the unification challenges of the DiSHA framework, allowing seamless adaptation across different LLM architectures such as LLaMA, RWKV, and Mistral but also enables it to significantly outperform LoRA in both computational efficiency and memory usage. In NLG tasks, Bone demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variant Pissa across the board. To verify the feasibility of Bone, we conducted experiments on two different LLM architectures (LLaMA2 Xu et al. (2023), RWKV6 Peng et al. (2024)). Based on the comparative experiments between Bone and Bat, it is evident that the non-linear updates of weight shards enhance the model’s performance ceiling.
@@
1. We propose a novel PEFT framework called DiSHA which significantly reduces the number of trainable parameters.
2. We proposed an efficient and user-friendly structure named Bone within a subset of the DiSHA design space and validated its effectiveness through extensive experiments.
3. We identified a potential issue in the DiSHA framework: the updates of weight shards are colinear. To address this, we improved upon Bone and proposed Bat, which transforms weight shard updates into non-linear ones, thereby enhancing the model’s performance ceiling.
4. Through extensive experiments, we validated the effectiveness of DiSHA, Bone, and Bat, achieving state-of-the-art (SOTA) performance.
@@ 2 Related Works
LoRA, as one of the most popular PEFT techniques, has demonstrated that the updates to the weights exhibit a low intrinsic rank during adaptation. In LoRA, for a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing the latter with a low-rank decomposition $W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$. During training, $W_{0}$ is frozen and does not receive gradient updates, while $A$ and $B$ contain trainable parameters. Note both $W_{0}$ and $\Delta W=BA$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise. For $h=W_{0}x$, our modified forward pass yields:
$h=W_{0}x+\Delta Wx=W_{0}x+BAx$ (1)
Since A and B are initialized with Gaussian noise and zeros in LoRA, the gradients can be very small, leading to slow convergence in the fine-tuning process. As a result, many LoRA variants have emerged, aiming to enhance LoRA’s capabilities by modifying the initialization or initial gradients, among other approaches.
LoRA+ extended this method by introducing independent learning rates for matrices $A$ and $B$ with a fixed ratio, improving the method’s efficiency. The DoRA Liu et al. (2024b) method combines weight decomposition to achieve learning capabilities similar to full fine-tuning without sacrificing LoRA’s inference efficiency. PiSSA optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction. Using Singular Value Decomposition (SVD), PiSSA initializes the dominant singular values and vectors to train these matrices, while keeping the residual matrix static during fine-tuning. OLoRA Büyükakyüz (2024) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal. This orthogonal initialization helps maintain the stability of the parameter space during optimization. LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $A$ and $B$ by computing the initial gradient, thereby closely approximating full fine-tuning.
The improvements brought by these LoRA variants validate that the updates to the weights exhibit a low intrinsic rank during adaptation and hold greater potential. However, they also introduce more complex initialization steps and increase preprocessing time. Therefore, we need to design novel structures to facilitate the exploration of a lower intrinsic rank.
@@ 3 Method
3.1 DiSHA: Dimension-Sharding Adaptation
Hypothesis Based on the conclusions of the experimental analysis of the gradient norm  9.1, we designed a novel framework called Dimension-Sharding Adaptation (DiSHA) to unlock lower intrinsic ranks and faster convergence by default. For a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing the latter with a low-rank decomposition $W=W_{0}+\Delta W=W_{0}+expand(D)$, where $D\in\mathbb{R}^{r_{1}\times r_{2}},expand(D)\in\mathbb{R}^{d\times k}$, and the rank $(r_{1},r_{2})\ll\min(d,k)$. During training, $W_{0}$ is frozen and does not receive gradient updates, while $D$ is a trainable parameter initialized to 0. Note both $W_{0}$ and $\Delta W=expand(D)$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise. For $y=W_{0}x$, where $x\in\mathbb{R}^{b\times l\times k}$, $y\in\mathbb{R}^{b\times l\times d}$, our modified forward pass yields:
$y=W_{0}x+\Delta Wx=W_{0}x+expand(D)x$ (2)
Here, we explain the concept of $expand(D)$. As shown in the DiSHA structure in Figure  3, we divide $W_{0}$ into multiple shards along the dimension of $D$. The $D$ matrix is then replicated according to the number of shards, and its dimensions are restored to match $W_{0}$. Finally, the updated shards are summed to update $W$.
3.2 DiSHA design space
In the previous section, we introduced the design principles of DiSHA and defined its trainable matrix $D\in\mathbb{R}^{r_{1}\times r_{2}}$. With DiSHA, you can customize the size of the trainable matrix $D$ based on the dimensions of the pre-trained weights. Since all shards of the pre-trained weights $W_{0}$ share a single $D$ for updates, as long as the expanded $expand(D)$ matches the dimensions of $W_{0}$, you can freely choose how to divide $W_{0}$ into shards. This flexibility makes the design space of DiSHA immense 2.
3.3 Bone: Block-Affine efficient update
In the previous section, we introduced the DiSHA framework, a method that significantly reduces the number of trainable parameters in the model. Due to the vast design space of the DiSHA framework, it is difficult for users to accurately determine suitable values for training. A good setting of $D\in\mathbb{R}^{r_{1}\times r_{2}}$ does not generalize well across all neural networks. For example, with the rise of technologies like GQA Ainslie et al. (2023) and MQA Shazeer (2019), the matrix dimensions of self-attention modules ($W_{k},W_{v}$) vary greatly among different LLM models. At the same time, computational efficiency and resource utilization also vary depending on the design space choices within the DiSHA framework.
To address these limitations, we proposed Block Affine Adaptation (Bone) based on the DiSHA framework. Inspired by LoRA, we modified $D\in\mathbb{R}^{r_{1}\times r_{2}}$ to $D\in\mathbb{R}^{r\times d}$ or $D\in\mathbb{R}^{r\times k}$. This allows the trainable parameter count to be controlled by a single parameter $r$, making it easy to adapt to any neural network. Additionally, we observed that $expand(D)x$ not only increases memory consumption but also reduces computational efficiency. From our analysis, when $D\in\mathbb{R}^{r\times d}$, $x\in\mathbb{R}^{b\times l\times k}$, $expand(D)\in\mathbb{R}^{d\times k}$, the operation $expand(D)x$ essentially performs a summation along the $k$-dimension, where $expand(D)$ involves replicating the necessary number of shards. Clearly, we can first $sum(x)$ along the $k$dimension and then compute it directly with $D$, where $sum(x)\in\mathbb{R}^{b\times l\times r}$. This approach not only reduces memory consumption but also improves computational efficiency. The update for $W$ is shown in the figure  3, and the Bone calculation formula is as follows:
$y=W_{0}x+\Delta Wx=W_{0}x+sum(x)D$ (3)
Figure 3: Diagram of LoRA, DiSHA, and Bone Architectures.
@@ 3.4 Bat: Block-Affine Transformation
We conducted extensive experiments on both NLU and NLG tasks to validate the effectiveness of Bone. It outperforms many LoRA variants and surpasses LoRA in terms of memory consumption and computational efficiency. However, Bone still has a potential issue, which stems from the shared update mechanism of the DiSHA framework. Specifically, all shards in the weights use the same trainable matrix for updates, causing the updates of all shards to be colinear. This restricts weight updates and, consequently, the model’s capabilities. To address the issue of linear correlation among shard updates, our initial idea was to use a trainable coefficient matrix to control the updates of different shards. However, this approach would increase additional parameters.
To avoid introducing additional parameters, we applied a linear transformation between the original weight $W_{0}$ shards and the trainable matrix, ensuring that the updates of different shards are nonlinearly related. Based on this, we proposed the Block Affine Transformation Adaptation (Bat) method. It extends Bone by further dividing both the trainable matrix and $W_{0}$ into smaller shards, where each shard interacts with the trainable matrix $D$ through a linear transformation, achieving nonlinear independence among shard updates. The specific formula is as follows:
$W_{k/r,d/r,r,r}$ $=Reshape(W_{d,k})$ (4)
$bone_{d/r,r,r}$ $=Reshape(bone_{r,d})$
$\Delta W_{d,k}$ $=Reshape(W_{k/r,d/r,r,r}bone_{d/r,r,r}+bone_{d/r,r,r})$
It is evident that Bat can be defined in many other ways, as illustrated in the figure 4. In Section 6, we compared the performance of different partitioning strategies for the structure. The formula for Bone-row is as follows:
$W_{d/r,k/r,r,r}$ $=Reshape(W_{d,k})$ (5)
$bone_{k/r,r,r}$ $=Reshape(bone_{r,k})$
$\Delta W_{d,k}$ $=Reshape(W_{d/r,k/r,r,r}bone_{k/r,r,r}+bone_{k/r,r,r})$
Figure 4: Diagram of Bat Structures in Different Design Spaces.